# Quasi-Periodic Oscillations and Parameter Constraints in ModMax Black Holes

### v1

@@ I Introduction
Nonlinear electrodynamics (NED) has emerged as an effective framework for generating regular and physically consistent field configurations in curved spacetime [70]. Unlike linear Maxwell theory, nonlinear electrodynamics characterised by gauge invariant Lagrangians dependent on the electromagnetic invariant $F=F_{\mu\nu}F^{\mu\nu}$ can yield stress-energy tensors exhibiting symmetries that imitate vacuum behaviour under radial boosts. Notably, among these theories, Born-Infeld (BI) electrodynamics has attracted considerable attention for its appearance in the low-energy regime of string theory. Furthermore, several other NED formulations [71, 72, 73, 74, 75] share desirable features with the Born-Infeld theory, such as finite electric field strength at the origin and finite total electrostatic energy. Another nonlinear theory worth mentioning in this regard is the Euler Heisenberg Nonlinear Electrodynamics (EH NLED) theory [76]. Both Euler Heisenberg and Born Infeld NLED preserve the electric-magnetic duality invariance. More recently a different non linear electrodynamics theory endowed with two symmetries was proposed [77]. This particular framework was characterised by the dimensionless parameter $\eta$ and for $\eta=0$ it reduces to the standard Maxwell theory. This modified version of electrodynamics, commonly referred to as the ModMax NLED theory, has stimulated extensive research in several directions, including the study of classical solutions [78] as well as supersymmetric extensions [79]. In what follows, we will review the features of the ModMax black hole adopted in this study.
@@
$\mathcal{I}=\frac{1}{16\pi}\int d^{4}x\sqrt{-g}\left(\mathcal{R}-4\mathcal{L}\right)$ (1)
here $\mathcal{R}$ is the Ricci scalar, $g$ denotes the determinant of the metric tensor $g_{\mu\nu}$ and $\mathcal{L}$ denotes the ModMax Lagrangian. This Lagrangian can also be written as [81]
$\mathcal{L}=\frac{1}{2}\left(\mathcal{S}\cosh{\eta}-\sqrt{\mathcal{S}^{2}+\mathcal{P}^{2}}\sinh{\eta}\right)$ (2)
where $\eta$ is a dimensionless quantity intrinsic to the ModMax theory. $\mathcal{S}$ and $\mathcal{P}$ are two real and pseudo-scalars respectively. The metric of a spherically symmetric charged ModMax black hole can be written as
@@ II.0.1 Equations of Motion
$\dot{r}^{2}=\mathcal{E}+g_{tt}\Big(1+\frac{{\mathcal{L}^{2}}}{r^{2}}\Big),$
$\dot{r}^{2}=\mathcal{E}+g_{tt}\Big(1+\frac{{\mathcal{L}^{2}}}{r^{2}}\Big).$ (8)
@@ V Monte Carlo Markov chain (MCMC) analysis
The posterior distributions suggest that the ModMax parameter $\eta$ remains of order unity across all black hole sources, implying a modest but consistent deviation from conventional electrodynamics. The relatively constrained values of $\eta$ and $Q/M$ across distinct mass scales indicate that nonlinear electromagnetic effects introduced by the ModMax framework could influence the quasi-periodic oscillation frequencies. Overall, the analysis reveals coherent signatures of the ModMax modification, reflecting its potential relevance in describing the dynamics of matter in strong-field regimes. We summarize the results in from the MCMC analysis in Table 2. We observe that the constraints on the mass $M$ is in good agreement with the observational values as tabulated in 1. The constrained values of $\eta$, $Q/M$ and $r/M$ are also shown in Table 2
@@
Finally, to connect our theoretical framework with observations, we employed a Markov Chain Monte Carlo (MCMC) analysis to constrain the parameters $\{M,\eta,Q/M,r/M\}$ of the ModMax black hole model using QPO data from six well-established black hole sources. These include four stellar-mass systems (GRO J1655–40, XTE J1550–564, GRS 1915+105, and H 1743–322), the intermediate-mass black hole M82 X-1, and the supermassive black hole Sgr A*, thereby covering a broad range of mass scales. We consider only the RP model in this case just for demonstration of the method. The results of the MCMC analysis are tabulated in Table 2, providing the constrained values of the above mentioned parameters. It is observed that the constrained mass value is in good agreement with the available observational data which is given in Table 1. The MCMC analysis also suggest that the ModMax parameter $\eta$ is of the order unity implying a modest but consistent deviation from conventional electrodynamics. It is worth noting that, although the observational data used in our study most likely correspond to rotating black holes, our analysis intentionally omits the spin parameter to isolate the intrinsic effects of nonlinear electrodynamics on QPO frequencies. The objective is to assess whether the nonlinear effect alone can reproduce the changes in frequency that are usually linked to rotation. Our findings suggest that a moderate nonlinear charge can partly mimic the influence of spin, allowing non-rotating black holes to show QPO features similar to those of rotating ones.

### v2

@@ I Introduction
Nonlinear electrodynamics (NED) has emerged as an effective framework for generating regular and physically consistent field configurations in curved spacetime [70]. Unlike linear Maxwell theory, nonlinear electrodynamics characterised by gauge invariant Lagrangians dependent on the electromagnetic invariant $F=F_{\mu\nu}F^{\mu\nu}$ can yield stress-energy tensors exhibiting symmetries that imitate vacuum behaviour under radial boosts. Notably, among these theories, Born-Infeld (BI) electrodynamics has attracted considerable attention for its appearance in the low-energy regime of string theory. Furthermore, several other NED formulations [71, 72, 73, 74, 75] share desirable features with the Born-Infeld theory, such as finite electric field strength at the origin and finite total electrostatic energy. Another nonlinear theory worth mentioning in this regard is the Euler Heisenberg Nonlinear Electrodynamics (EH NLED) theory [76]. Both Euler Heisenberg and Born Infeld NLED preserve the electric-magnetic duality invariance. More recently a different non linear electrodynamics theory endowed with two symmetries was proposed [77]. This particular framework was characterised by the dimensionless parameter $\eta$ and for $\eta=0$ it reduces to the standard Maxwell theory. This modified version of electrodynamics, commonly referred to as the ModMax NLED theory, has stimulated extensive research in several directions, including the study of classical solutions [78] as well as supersymmetric extensions [79]. In what follows, we will review the features of the ModMax black hole adopted in this study. It is important to emphasize that NED effects, including those described by ModMax theory, are far too weak to be experienced in everyday or laboratory scale electromagnetic phenomena. Such nonlinear corrections become significant only in the presence of extremely strong electromagnetic and/or gravitational fields, as encountered in the vicinity of black holes. In this sense, black holes serve as natural laboratories for probing and testing nonlinear electrodynamic theories.
@@
$\mathcal{I}=\frac{1}{16\pi}\int d^{4}x\sqrt{-g}\left(\mathcal{R}-4\mathcal{L}^{\prime}\right)$ (1)
here $\mathcal{R}$ is the Ricci scalar, $g$ denotes the determinant of the metric tensor $g_{\mu\nu}$ and $\mathcal{L}^{\prime}$ denotes the ModMax Lagrangian. This Lagrangian can also be written as [81]
$\mathcal{L}^{\prime}=\frac{1}{2}\left(\mathcal{S}\cosh{\eta}-\sqrt{\mathcal{S}^{2}+\mathcal{P}^{2}}\sinh{\eta}\right)$ (2)
where $\eta$ is a dimensionless quantity intrinsic to the ModMax theory. $\eta$ can be interpreted as an effective coupling parameter that characterizes the strength of nonlinear deviations from standard Maxwell electrodynamics. $\mathcal{S}$ and $\mathcal{P}$ are two real and pseudo-scalars respectively. The metric of a spherically symmetric charged ModMax black hole can be written as
@@ II.0.1 Equations of Motion
$\dot{r}^{2}=\mathcal{E}^{2}+g_{tt}\Big(1+\frac{{\mathcal{L}^{2}}}{r^{2}}\Big)$ (8)
@@ V Monte Carlo Markov chain (MCMC) analysis
The posterior distributions suggest that the ModMax parameter $\eta$ remains of order unity across all black hole sources, implying a modest but consistent deviation from conventional electrodynamics. The relatively constrained values of $\eta$ and $Q/M$ across distinct mass scales indicate that nonlinear electromagnetic effects introduced by the ModMax framework could influence the quasi-periodic oscillation frequencies. The relatively large $Q/M$ ratios inferred in this analysis are difficult to reconcile with standard astrophysical expectations, since any macroscopic charge on a black hole would be rapidly neutralized by the surrounding plasma. We therefore suggest that the charge parameter in the ModMax framework should be interpreted as an effective parameter, rather than as a literal astrophysical electromagnetic charge. From this perspective, the inferred charge may instead encode additional degrees of freedom arising from nonlinear electrodynamic effects.
We summarize the results in from the MCMC analysis in Table 2. We observe that the constraints on the mass $M$ is in good agreement with the observational values as tabulated in 1. The constrained values of $\eta$, $Q/M$ and $r/M$ are also shown in Table 2
@@
Finally, to connect our theoretical framework with observations, we employed a Markov Chain Monte Carlo (MCMC) analysis to constrain the parameters $\{M,\eta,Q/M,r/M\}$ of the ModMax black hole model using QPO data from six well-established black hole sources. These include four stellar-mass systems (GRO J1655–40, XTE J1550–564, GRS 1915+105, and H 1743–322), the intermediate-mass black hole M82 X-1, and the supermassive black hole Sgr A*, thereby covering a broad range of mass scales. In this work, we restrict our analysis to the RP model solely for the purpose of demonstrating the methodology. Consequently, the parameter constraints obtained and the physical interpretations inferred from the MCMC analysis are specific to the RP model considered here. Different QPO models may lead to quantitatively different constraints and potentially distinct physical interpretations, and exploring such model dependence is left for future investigations.
The results of the MCMC analysis are tabulated in Table 2, providing the constrained values of the above mentioned parameters. It is observed that the constrained mass value is in good agreement with the available observational data which is given in Table 1. The MCMC analysis also suggest that the ModMax parameter $\eta$ is of the order unity implying a modest but consistent deviation from conventional electrodynamics.
It is worth noting that, although the observational data used in our study most likely correspond to rotating black holes, our analysis intentionally omits the spin parameter to isolate the intrinsic effects of nonlinear electrodynamics on QPO frequencies. However, the deliberate omission of rotation inevitably introduces a degeneracy in the parameter estimation. In particular, the inferred constraints on the nonlinear electrodynamic parameter $\eta$ may effectively absorb frequency shifts that would otherwise be attributed to black hole spin. As a result, although the primary objective of this manuscript is to examine whether nonlinear electrodynamic effects alone can reproduce frequency variations commonly associated with rotation, it should be emphasized that the best-fit values of the NED parameter obtained from the MCMC analysis are unlikely to arise purely from nonlinear effects. Instead, they should be interpreted as effective parameters that may implicitly encode, at least in part, the influence of rotational dynamics. Our findings suggest that a moderate nonlinear charge can phenomenologically accommodate the observed QPO frequencies within the present model framework.
Next, we compare our results obtained within the RP model with those of Ref. [105], in which a different nonlinear electrodynamic model was investigated in the context of observed QPOs. While the two studies differ in the underlying NLED formulation, a meaningful comparison can be made at the level of inferred parameter trends and effective physical interpretations. In our ModMax-based analysis, the nonlinear electrodynamic effects are characterized by the parameter $\eta$, whereas in the comparative study the corresponding role is played by the parameter $b$. Despite this difference in parametrization, both analyses yield broadly consistent mass estimates for the considered sources, which are in good agreement with independent astrophysical constraints. A comparison of the $r/M$, reveals that both models predict the value of $r/M$ within $r/M\sim 5$1212. While the exact numerical values differ between the two studies, the overall scale and ordering across sources remain comparable. In both models, the inferred values of the charge-to-mass ratio $Q/M$ are found to be close to unity. Although the corresponding $Q/M$ values obtained in Ref. [105] are somewhat lower, they remain of the same order of magnitude. Such large values are difficult to interpret as a literal astrophysical electric charge, since any real charge would be rapidly neutralized in an astrophysical environment. This supports the interpretation of $Q/M$ as an effective parameter that captures the influence of nonlinear electrodynamic effects, rather than representing a true electromagnetic charge of the black hole. Similarly, although both $\eta$ and $b$ act as measures of deviation from the Maxwell limit, their posterior distributions are not directly comparable on a one-to-one basis. Instead, they should be viewed as model-specific parameters whose numerical values depend on the particular functional form of the nonlinear electrodynamic Lagrangian. Despite these differences, both parameters consistently take nonzero values greater than unity across multiple sources, suggesting that nonlinear electromagnetic effects may play a phenomenologically relevant role in shaping QPO frequencies within the RP model. Overall, this comparison highlights that certain inferred quantities, such as the black hole mass and the characteristic QPO radius, remain relatively robust across different nonlinear electrodynamic models. The constraints on the effective charge-to-mass ratio are found to be of order unity in both models, although slightly lower values are obtained in Ref. [105]. In contrast, while the nonlinear electrodynamic parameters in the two models are of the same order of magnitude, $\mathcal{O}(10^{0})$, their numerical values differ significantly. This model dependence shows the importance of exercising caution when attributing direct physical meaning to individual parameter values and motivates future studies exploring multiple QPO prescriptions and nonlinear electrodynamic theories within a unified statistical framework.